OpenStudy (geekfromthefutur):

Maria is comparing the prices of two window cleaning companies. Company A charges $6 per window and an additional $12 as service charges. Company B charges $5 per window and an additional $15 as service charges.

OpenStudy (anonymous):

I jst helped u

OpenStudy (geekfromthefutur):

no ik sorry my internet gets slow when theres so many comments

OpenStudy (anonymous):

what are you trying to figure out?

OpenStudy (geekfromthefutur):

Part A: Write equations to represent Company A and Company B's total charges for cleaning a certain number of windows. Define the variable used in the equations. (4 points) Part B: Which company would charge less for cleaning 8 windows? Justify your answer. (3 points) Part C: How much money is saved by using the services of Company B instead of Company A to clean 6 windows? (3 points)

OpenStudy (anonymous):

A: 12+6x B: 15+5x x= amt of windows cleaned

OpenStudy (anonymous):

I understand @geekfromthefutur

OpenStudy (anonymous):

A: 12+48= $60 B: 15+40= $55 Company B is cheaper

OpenStudy (anonymous):

A: 12+36= $48 B: 15+30= $45 you save $3

OpenStudy (anonymous):

there you go... hope this helps

OpenStudy (geekfromthefutur):

thank you (: can u help me with a few more

OpenStudy (anonymous):

sure

OpenStudy (geekfromthefutur):

ok thank you :3

OpenStudy (geekfromthefutur):

Emily uses her college photocopier machine to photocopy at the rate of $0.08 per page. She decides to rent a photocopier machine for $90 a year. The cost of photocopying using the rented machine is $0.03 per page. Part A: Write an inequality that can be used to calculate the number of pages that Emily should photocopy in a year so that the amount she pays for the rented machine is less than the college machine. Define the variable used. (5 points) Part B: How many pages should Emily print in a year to justify renting the photocopier? Show your work. (5 points)

OpenStudy (anonymous):

90+0.03x < 0.08x part B may take a while

OpenStudy (geekfromthefutur):

ok take your time (:

OpenStudy (anonymous):

mkk

OpenStudy (geekfromthefutur):

u still there @elementwielder

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

90+0.03(1800)=0.08(1800)

OpenStudy (geekfromthefutur):

oh ok is that part B ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

simplified that would be 144=144

OpenStudy (geekfromthefutur):

oh ok thanks dude your the best this math i freken suck at this lol

OpenStudy (anonymous):

no i suck... this is just easier for me

OpenStudy (anonymous):

and your welcome anytime

OpenStudy (geekfromthefutur):

ok so its 90+0.03(1800)=0.08(1800) and i then say simplified that would be 144=144 right?

OpenStudy (anonymous):

ya

OpenStudy (geekfromthefutur):

oh thx bro hey can u help me with 2 more im really sorry iv im wasting any of your time >.<

OpenStudy (anonymous):

nah i have plenty of time...

OpenStudy (geekfromthefutur):

oh thx bro (;

OpenStudy (anonymous):

your welcome...

OpenStudy (geekfromthefutur):

In a certain town the temperature, x in degrees Celsius on a certain day is described by two statements: If 3 times the temperature is increased by 2, the temperature is still less than 14°C. Twice the temperature minus 7 is greater than -11°C. Part A: Create a compound inequality to represent the temperature range. (3 points) Part B: Can the temperature in this town be 5°C? Justify your answer by solving the inequalities in Part A. (3 points) Part C: The average temperature in another town is 3°C but the actual temperature is within 4°C of the average. Write and solve an inequality to find the range of temperature in this town. (4 points)

OpenStudy (geekfromthefutur):

Part A: Solve C = 5 over 9 (F - 32) for F. (4 points) Part B: Determine the value of F when C = 8 °C. (2 points) Part C: Solve –np – 80 > 60 for n. Show your work. (4 points)

OpenStudy (anonymous):

-11 < 2x-7 or 3x+2 < 14

OpenStudy (anonymous):

NO... 3(5)+2>14 simplified: 17>14

OpenStudy (anonymous):

-1 is less than or equal to X is less than or equal to 7

OpenStudy (anonymous):

those 3 above are four the 1st section of questions

OpenStudy (geekfromthefutur):

oh ok thanks bro :3

OpenStudy (anonymous):

\[C =\frac{ 5 }{ 9\times(F -32) }\]

OpenStudy (anonymous):

multiply F to both sides solve right side: \[\frac{ 5 }{ 288 }\] divide C to both sides... \[F=\frac{ 5 }{ -288C }\]

OpenStudy (anonymous):

you can solve B from there... im not sure how to do C

OpenStudy (geekfromthefutur):

oh ok so the first question u said C= 5 over 9 x (F -32) is for part A right jw

OpenStudy (anonymous):

ya under your last post the following 2 of my posts were for part A

OpenStudy (geekfromthefutur):

oh ok thx

OpenStudy (geekfromthefutur):

(;

OpenStudy (anonymous):

your welcome sorry i couldnt help you out with Part C

OpenStudy (geekfromthefutur):

its cool thanks tho (:

OpenStudy (anonymous):

your welcome...

OpenStudy (geekfromthefutur):

hey i was wondering how do i divide C from both sides?

OpenStudy (anonymous):

for?

OpenStudy (geekfromthefutur):

for part B i was trying to figure out how 2 divide C from both sides

OpenStudy (anonymous):

well i have that set up a few messages up... so you dont need to do the hard work

OpenStudy (geekfromthefutur):

oh ok thx :3

OpenStudy (anonymous):

your welcome... lol

OpenStudy (geekfromthefutur):

oh ok thts all the questions thank you so much for all the help :)

OpenStudy (anonymous):

your welcome anytime im on or not on just ask and ill answer asap

OpenStudy (geekfromthefutur):

oh ok thx :3

OpenStudy (anonymous):

welcome